OpenStudy (abmon98):

find dy/dx of sinx+2cosy=1

OpenStudy (anonymous):

=cosx-2siny dy/dx=0 =>cosx=2siny dy/dx =>dy/dx=cosx/2siny

OpenStudy (abmon98):

thank you @zaibali.qasmi could you help me in one more question e^xy=4

OpenStudy (anonymous):

what have to do it...???derivatie..?

OpenStudy (abmon98):

yes.

OpenStudy (anonymous):

thenn see it.. e^xy*(1.y+xdy/dx)=0 =>ye^xy+xe^xydy/dx =>ye^xy=-xe^xydy/dx =>dy/dx=-y/x